# Life Sequence Transformer: Generative Modelling for Counterfactual Simulation

### v1

@@ Appendix D Experimental details
The third exercise examines whether the model can replicate age-based differences in the duration of unemployment benefits, specifically the mobility allowance, a now-discontinued income support program for workers displaced by firm closures in Italy. As shown in Figure 5 less than 1% of records in our global sample fall under this category.
@@
Under the policy rules in place until January 2017, the maximum benefit duration hinged on the worker’s age at the time of job loss. Workers over 40 were eligible for up to 24 months of benefits, while those under 40 could receive income support for 12 months [Brunello and Miniaci, 1997]. This sharp policy threshold makes it an ideal setting to test whether the model internalises and reflects such causal discontinuities in allowance’s duration.
@@
We drew individuals with at least one registered episode of mobility allowance from the test set, who also have only standard employment or retirement spells in their histories. To increase homogeneity, we exclude individuals with any history of self-employment, artistic or athletic careers, or voucher-based contracts. Additionally, we retain only those employed in northern Italian provinces at the time of firm closure; a worker from the south is allowed one additional year of benefits.
We restrict the sample to workers who began receiving the benefit between the ages of 30 and 50, and who have less than 1% missing data on key variables: economic sector (ATECO), work province, income, and job title. The resulting sample consists of 1502 individuals.
@@
We provide the model with each individual’s life trajectory up to the year we first observe the unemployment episode. The model then generates the sequence from that point onward. This setup allows us to test whether the simulated unemployment duration reflects the correct age-based discontinuity in benefit length.
Out-of-sample Generation
@@
The purpose of this experiment is to evaluate the structural coherence of out-of-sample generations over an extended horizon. For each generated sequence, we identify the first year in which the model produces a structural inconsistency. Conistency means that each calendar year must begin/end with an End-of-Year <EOY> token, followed by a valid combination of monthly tokens and corresponding durations. These can be empty, i.e. no event, or include a valid life event, but must respect the order: month, then optional content, then duration. Inside each year, there must be at least one event, even if empty (so no repeated <EOY> tokens). A duration token must be followed either by another month token or an <EOY> token indicating the start of the next year, and <EOY> tokens must always be followed by a month token.
@@ Appendix E Additional results
Table 6: Decoder-only test set metrics for next-token prediction, macro-averaged across the vocabulary, conditioning on 0, 5, 1 and 10 initial years of sequence.
Accuracy Precision Recall F1 Score Perplexity
0 known years 0.6225 0.6868 0.6225 0.6427 0.7288
1 known year 0.6773 0.7216 0.6773 0.6895 0.6937
5 known years 0.6883 0.7285 0.6883 0.6991 0.6727
10 known years 0.7039 0.7409 0.7039 0.713 0.6466
Figure 8: Real and decoder-only model-generated (blue) average pension ages—measured in months—for male individuals born between 1940 and 1950, stratified by birth month (January = M1, December = M12). Each point indicates the cohort-specific average retirement age for individuals born in January or December, with vertical bars representing a one standard deviation range.
Figure 9: Real and encoder-decoder model-generated trajectories of average monthly income for working mothers in a 10-year window before and after their first maternity leave episode (Years From Maternity = 0). The solid lines represent the average monthly income, while the shaded areas denote the one standard deviation range.
Figure 10: ATECO grouped real and decoder-only model-generated trajectories of average monthly income for working mothers in a 10-year window before and after their first maternity leave episode (Years From Maternity = 0). The solid lines represent the average monthly income, while the shaded areas denote the one standard deviation range.
Figure 11: ATECO grouped real and encoder-decoder model-generated trajectories of average monthly income for working mothers in a 10-year window before and after their first maternity leave episode (Years From Maternity = 0). The solid lines represent the average monthly income, while the shaded areas denote the one standard deviation range.
@@
Figure 12: Real and decoder-only model simulated average duration with 90% confidence interval of mobility allowance receipt duration as a function of worker age at job loss in months in a 10-year bandwidth around the 40 years old threshold
@@
Figure 13: Density plot for the decoder-only out-of-sample year when the model generates incorrect sequences, 5% of the observations were able to complete the 20 years of generation without failures.
Figure 14: Density plot for the encoder-decoder out-of-sample year when the model generates incorrect sequences, 0% of the observations were able to complete the 20 years of generation without failures.
@@
Figure 15: Two-dimensional projection of the learned embedding space from the encoder-decoder model, visualized using the PaCMAP dimensionality reduction technique [Wang et al., 2021]. Each point represents a token from the vocabulary ($\lvert\mathcal{V}\rvert=661$), with different colours denoting different token categories. Several regions of interest are highlighted with zoomed-in views, each showing the top three nearest neighbours of a specific token. Neighbour relationships are computed using cosine distance in the original high-dimensional embedding space ($d=240$).
Figure 16: Two-dimensional projection of the learned embedding space from the decoder-only model, visualized using the PaCMAP dimensionality reduction technique [Wang et al., 2021]. Each point represents a token from the vocabulary ($\lvert\mathcal{V}\rvert=661$), with different colours denoting different token categories. Several regions of interest are highlighted with zoomed-in views, each showing the top three nearest neighbours of a specific token. Neighbour relationships are computed using cosine distance in the original high-dimensional embedding space ($d=240$).

### v2

@@ D.1 Mobility allowance duration
The first exercise examines whether the model can replicate age-based differences in the duration of unemployment benefits, specifically the mobility allowance, a now-discontinued income support program for workers displaced due to firm closures in Italy. As shown in Figure 6, less than 1% of records in our global sample fall under this category.
@@
Under the policy rules in place until January 2017, the maximum benefit duration depended on the worker’s age at the time of job loss. Workers over 40 were eligible for up to 24 months of benefits, while those under 40 could receive income support for up to 12 months (Brunello and Miniaci, 1997). This sharp policy threshold makes it an ideal setting to test whether the model internalises and reflects such causal discontinuities in allowance duration.
@@
We drew individuals with at least one registered episode of mobility allowance from the test set who also have only standard employment or retirement spells in their histories. To increase homogeneity, we exclude individuals with any episode of self-employment, artistic or athletic careers, or voucher-based contracts. Additionally, we retain only those employed in northern Italian provinces at the time of firm closure, since a worker from the south was allowed one additional year of benefits after 40 years of age. We restrict the sample to workers who began receiving the benefit between the ages of 30 and 50, and who have less than 1% missing data on key variables: economic sector (ATECO), work province, income, and job title. The resulting sample consists of 1,502 individuals.
@@
For each individual, we provide the model with their life trajectory up to the year we first observe the receipt of unemployment benefits. The model then generates the sequence from that point onward. This setup allows us to test whether the simulated unemployment subsidy duration reflects the correct age-based discontinuity.
To quantify the effect, we exploit the policy cutoff at 40 years of age (480 months) and estimate a local average treatment effect (ATE) within symmetric windows $h\in\{12,48,96\}$ months around the threshold. Let $A_{i}$ denote age at first unemployment, and define the sample and treatment within the window, and treatment dummy $T_{i}$
$\mathcal{I}(h)=\{i:|A_{i}-480|\leq h\},\qquad T_{i}=\mathbb{I}[A_{i}\geq 480].$
Because an exogenous age cutoff determines the treatment, differences in unemployment duration just above and just below the threshold can be interpreted as a causal effect of the policy rule. Accordingly, the treatment effect is defined as a local difference in means, directly comparable between the empirical data and the model-generated sequences
$\widehat{\text{ATE}}(h)=\mathbb{E}[Y_{i}\mid T_{i}{=}1,\,i\in\mathcal{I}(h)]-\mathbb{E}[Y_{i}\mid T_{i}{=}0,\,i\in\mathcal{I}(h)],$
where $Y_{i}$ is the observed duration of the first unemployment spell. This corresponds to a simple difference in means above vs. below the cutoff inside the window, with bootstrap confidence intervals obtained by resampling individuals from $\mathcal{I}(h)$, and the signed discrepancy dependent on the window
$\Delta(h)=\widehat{\text{ATE}}_{\text{model}}(h)-\widehat{\text{ATE}}_{\text{emp}}(h).$
D.2 Child Penalty
The second exercise evaluates the model’s capacity to reproduce the income trajectory experienced by women following childbirth, a well-established finding in the causal literature on gender inequality in the labour market (Kleven et al., 2019; Casarico and Lattanzio, 2023).
From the test set, we selected a sample of women who had only held standard employment contracts throughout their recorded work histories. We identify mothers as those who experienced at least one maternity leave spell, defined by a maternity leave of at least two months in a given year. To ensure sample homogeneity, we exclude women who were ever classified as managers or professional athletes, as these careers may follow atypical trajectories. We retain only those with less than 1% missing data across key variables (ATECO sector, work province, income, and job title). Finally, we restrict to women whose first childbirth occurs between the 5th and 20th observed year, ensuring sufficient pre- and post-treatment periods, and only mothers working in the southern regions of Italy. The resulting evaluation set contains 631 mothers.
For each mother, we truncate the observed sequence at maternity leave ($t=0$) and generate forward from that point. In addition, we construct offset-1 variants where the sequence is truncated one year before the maternity leave (t=1t=-1). In these offset-1 runs, the model is not forced to generate a maternity event in the sequence, providing a natural basis for counterfactual comparison.
Unlike pension and unemployment, where treatment is determined by exogenous or quasi-exogenous rules, fertility is a choice-driven event. To address this treatment endogeneity, we adopt an event study design that compares sharp changes in labour market outcomes around the birth of the first child for mothers (M) relative to non-mothers (NM). The timing of childbirth creates a discrete shock to mothers’ careers, while other determinants of labour market outcomes would be expected to evolve smoothly over time. This approach also allows us to trace the full dynamic trajectory of the effects with precision, exploiting individual-level variation in the timing of first births (Kleven et al., 2019).
For individual $i$ in year $s$ with event time $t$ (measured in years since first maternity leave) and treatment group $T\in\{\text{M, NM}\}$, we run the following regression separately for mothers and non-mothers:
$Y^{T}_{ist}=\sum_{j\neq-1}\alpha^{T}_{j}\mathbf{1}[t=j]\;+\;\sum_{k}\beta^{T}_{k}\mathbf{1}[\text{age}=k]\;+\;\sum_{y}\gamma^{T}_{y}\mathbf{1}[\text{year}=y]\;+\;\nu^{T}_{ist},$
omitting t=1t=-1 as the reference period. This yields dynamic effects $\hat{\alpha}^{T}_{t}$ relative to the year before birth, while the flexible age and calendar-year fixed effects control for lifecycle and macroeconomic trends.
To compute percentage penalties, we normalise with the counterfactual income path absent childbirth, defined by the predicted values from the age and year components,
$\widetilde{Y}^{T}_{ist}=\sum_{k}\hat{\beta}^{T}_{k}\mathbf{1}[\text{age}=k]+\sum_{y}\hat{\gamma}_{y}\mathbf{1}[\text{year}=y],\qquad\widehat{P}^{T}_{t}=\frac{\hat{\alpha}^{T}_{t}}{\mathbb{E}[\widetilde{Y}^{T}_{ist}\mid t]}.$
We construct two types of controls, i.e., non-mothers (NM), against which the post-birth path of mothers is compared:
  1. 1.
    Matched non-mothers. Using propensity scores based on pre-birth characteristics that correlated both with treatment assignment and outcomes (age, year of birth, area of birth, year start working, first sector), we match each mother to a non-mother on common support. The matched control is aligned to the mother’s event time so that $t=0$ corresponds to the mother’s first birth year.
  2. 2.
    Simulated no-maternity trajectories. For each true mother, we use the offset-1 simulations where the model does not generate any maternity leave episode. We then average them per mother-year to form a synthetic counterfactual trajectory aligned to the event time of the true.
At event time $t$, we take the event-study outputs for mothers (treated, $\text{T}=\text{M}$) and the chosen control group (non-mothers or simulated no-maternity paths, $\text{T}=\text{NM}$) and define the percentage child penalty as the difference in the estimated event coefficients
$P_{t}\;=\;\widehat{P}^{(\text{T}=\text{M})}_{t}\;-\;\widehat{P}^{(\text{T}=\text{NM})}_{t}$, (2)
where each $\widehat{P}^{(\cdot)}_{t}$ is already reported in percentage units by the event-time regression. We summarise the dynamic path by averaging over an event-time window
$P_{a\text{--}b}\;=\;\frac{1}{b-a+1}\sum_{t=a}^{b}P_{t}.$ (3)
For each pairing (empirical matching and model-based control) and each event time $t$ used in the window, we observe the estimated percentage effects and their standard errors for both arms:
$\Big(\widehat{P}^{(\text{T}=\text{M})}_{t},\,\widehat{\sigma}^{(\text{T}=\text{M})}_{t}\Big)\quad\text{and}\quad\Big(\widehat{P}^{(\text{T}=\text{NM})}_{t},\,\widehat{\sigma}^{(\text{T}=\text{NM})}_{t}\Big)$.
We then implement a parametric (Gaussian) bootstrap:
  1. 1.
    For every $t\in\{a,\ldots,b\}$, draw independently
    $\widetilde{P}^{(\text{T}=\text{M})}_{t}\sim\mathcal{N}\!\big(\widehat{P}^{(\text{T}=\text{M})}_{t},\,(\widehat{\sigma}^{(\text{T}=\text{M})}_{t})^{2}\big),\qquad\widetilde{P}^{(\text{T}=\text{NM})}_{t}\sim\mathcal{N}\!\big(\widehat{P}^{(\text{T}=\text{NM})}_{t},\,(\widehat{\sigma}^{(\text{T}=\text{NM})}_{t})^{2}\big).$ (4)
  2. 2.
    Form the bootstrap penalty path and its window average:
    $\widetilde{P}_{t}\;=\;\widetilde{P}^{(\text{T}=\text{M})}_{t}\;-\;\widetilde{P}^{(\text{T}=\text{NM})}_{t},\qquad\widetilde{P}_{a\text{--}b}\;=\;\frac{1}{b-a+1}\sum_{t=a}^{b}\widetilde{P}_{t}.$ (5)
  3. 3.
    Repeat the previous steps $B$ times. The $95\%$ confidence interval for $P_{a\text{--}b}$ is the percentile interval over these runs.
Let $P^{\text{model}}_{a\text{--}b}$ and $P^{\text{emp}}_{a\text{--}b}$ denote the window-averaged penalties computed from the model-based control and from the empirical matched control, respectively. We report the signed difference and its absolute value:
$\Delta_{a\text{--}b}\;=\;P^{\text{model}}_{a\text{--}b}-P^{\text{emp}}_{a\text{--}b}.$ (6)
A confidence interval for $\Delta_{a\text{--}b}$ is obtained from paired parametric draws: on each bootstrap iteration $b$, we simulate $\widetilde{P}^{\text{model}}_{a\text{--}b}$ and $\widetilde{P}^{\text{emp}}_{a\text{--}b}$ from their respective Gaussians (using the reported standard errors at each $t$) and record $\widetilde{\Delta}_{a\text{--}b}=\widetilde{P}^{\text{model}}_{a\text{--}b}-\widetilde{P}^{\text{emp}}_{a\text{--}b}$. The percentile interval over these runs is reported as the $95\%$ confidence interval.
D.3 Retirement Timing
As a final causal benchmarking exercise, we evaluate the model’s ability to replicate the well-documented finding that the month of birth influences retirement timing (Ardito et al., 2020). Specifically, individuals born earlier in the calendar year tend to retire slightly later, due to institutional and lifecycle timing mechanisms.
We select from the test set all male individuals born between 1940 and 1950, restricting the sample to those born in either January or December. To ensure a clean employment trajectory, we include only those who have been continuously and exclusively employed in standard employment, excluding individuals who were ever self-employed, para-subordinate, unemployed, or employed under special regimes (e.g., artists, athletes, or voucher-based contracts). We further restrict the sample to those whose first pension spell occurs between ages 55 and 68, and who exhibit no long gaps in their administrative records. To reduce noise from incomplete information, we also exclude individuals with missing values in key attributes such as work province, economic sector (ATECO), income, and job title. After applying these criteria, the final evaluation set comprises 1,010 individuals.
For each of these individuals, we truncate their observed sequence prior to retirement and let the model generate forward until 2015. We consider two cutoff horizons: one year before the observed retirement (1y) and four years before retirement (4y), with $K=8$ simulations per individual. The longer horizon increases predictive uncertainty but avoids anchoring the model too closely to the retirement event.
The empirical average treatment effect is the difference in mean retirement age (in months) between January- and December-born individuals in the evaluation set:
$\widehat{\text{ATE}}\;=\;\mathbb{E}[Y_{i}\mid T_{i}{=}1]\;-\;\mathbb{E}[Y_{i}\mid T_{i}{=}0],$
where $T_{i}=1$ if born in January and $T_{i}=0$ if born in December. In the pension application, treatment assignment is effectively exogenous, since the month of birth is orthogonal to individual career choices and labour market conditions. As a result, the treatment effect can be interpreted as a clean difference in means between the two groups of retirees.
D.4 Out-of-sample Generation
@@
The purpose of this experiment is to evaluate the structural coherence of out-of-sample generations over an extended horizon. For each generated sequence, we identify the first year in which the model produces a structural inconsistency. Consistency means that each calendar year must begin/end with an End-of-Year <EOY> token, followed by a valid combination of monthly tokens and corresponding durations. These can be empty, i.e., no event, or include a valid life event, but must respect the order: month, then optional content, then duration. For each year, there must be at least one event, even if empty (so no repeated <EOY> tokens). A duration token must be followed either by another month token or an <EOY> token indicating the start of the next year, and <EOY> tokens must always be followed by a month token.
@@ Appendix E Additional results
Figure 9 compares the real and simulated average duration of mobility allowance receipt as a function of age at displacement. Taken together, these results suggest that while the model does not reproduce the sharp discontinuity in unemployment benefit duration at the exact policy cutoff, it does capture the broader age gradient and approximates the average magnitude of the causal effect when evaluated in wider horizons around the treatment. This underscores both a limitation, i.e., the difficulty of learning from rare, rule-driven discontinuities, and a strength, i.e., when sufficient natural variation is present, the model’s simulated treatment effects align closely with their empirical counterparts.
@@
Figure 9: Real and model simulated average duration with 90% confidence interval of mobility allowance duration as a function of worker age at job loss in months in a 10-year bandwidth around the 40-year-old threshold.
Figure 10 provides a first look at how mothers’ average earnings trajectories compare to two types of counterfactual controls: model-generated synthetic paths without maternity events (left) and propensity-score–matched non-mothers (right). Both approaches highlight the earnings divergence that emerges around childbirth, but they differ in persistence: the simulated control yields a child penalty that appears to persist over time, whereas with propensity-score matching, the point estimates attenuate toward zero by year 10 after birth. Notably, in the two years before maternity, we already observe a slight increase followed by a decrease in earnings, even before childbirth occurs, highlighting how difficult it is to pinpoint when treatment truly begins and what baseline trajectory should serve as the counterfactual. This contrast raises the broader question of what constitutes a good counterfactual group and appropriate conceptual questions: at what point does a woman become a mother for the purpose of defining counterfactuals, and when are the relevant choices actually made, surely not at the precise year the first child arrives?
Figure 10: Average annual earnings of mothers (circle dots) compared to two types of controls (triangle dots). The left panel uses model-generated controls based on offset–1 simulations without maternity events, and the right panel uses propensity-score–matched non-mothers. Shaded areas denote one standard deviation.
Figure 11 then summarises the event time coefficients estimated as a percentage of the counterfactual outcome absence childbirth, i.e., $\widehat{P}^{T}_{t}$, relative to the year before the maternity event (t=1t=-1). The model-generated simulated controls (left panel) and propensity-score–matched non-mothers (right panel) yield broadly similar patterns, though with notable differences. In the simulated control case, the immediate short-run penalty tends to be smaller, but the estimates suggest a more persistent medium- to long-run effect. By contrast, the propensity-score comparison shows a sharper initial earnings drop and a quicker recovery. Taken together, these results show that the model reproduces empirical child penalties quite closely across horizons, while also revealing how alternative counterfactual constructions can shift the interpretation of the penalty’s timing and persistence.
Figure 11: Event time coefficients estimated as a percentage of the counterfactual outcome $\widehat{P}^{T}_{t}$ relative the year before the maternity event (t=1t=-1). The left panel uses model-generated synthetic controls based on offset–1 simulations without maternity events, while the right panel uses propensity-score–matched non-mothers as controls. The horizontal axis shows event time $t$ (years since birth), and vertical bars denote 95% confidence intervals.
Figure 12: ATECO grouped real and model-generated trajectories of average annual earnings for working mothers in a 10-year window before and after their first maternity leave episode (Years From Maternity = 0). The solid lines represent the average monthly income, while the shaded areas denote the one standard deviation range.
Figure 13 illustrates how the estimated retirement age varies with the choice of conditioning horizon. When the model is conditioned four years before retirement (left panel), it produces a clearer discontinuity between January- and December-born cohorts, whereas conditioning only one year before retirement (right panel) yields much smaller differences. This contrast underscores how sensitive the results are to where the observed sequence is cut: cutting further away forces the model to generate more of the trajectory, which amplifies the birth-month effect, while cutting closer reduces this reliance on simulated paths. These patterns highlight not only the model’s ability to reproduce well-documented empirical regularities but also the conceptual challenge of deciding how much information to condition on when constructing counterfactuals in retirement settings.
Figure 13: Real and model-generated (blue) average retirement ages, measured in months, for male individuals born between 1940 and 1950. The left panel shows results using a cutoff four years before the observed retirement year, while the right panel shows results using a cutoff one year before. Each point indicates the cohort-specific average retirement age for January (M1) and December (M12), with vertical bars representing one standard deviation.
@@
Figure 14: Density plot for the out-of-sample year when the model generates incorrect sequences, 5% of the observations were able to complete the 20 years of generation without failures.
@@
Figure 15: Two-dimensional projection of the learned embedding space from the model, visualised using the PaCMAP dimensionality reduction technique (Wang et al., 2021). Each point represents a token from the vocabulary ($\lvert\mathcal{V}\rvert=661$), with different colours denoting different token categories. Several regions of interest are highlighted with zoomed-in views, each showing the top three nearest neighbours of a specific token. Neighbour relationships are computed using cosine distance in the original high-dimensional embedding space ($d=240$).